OpenStudy (anonymous):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____.

OpenStudy (gorv):

y=6-5x put it in 2nd Equation 5x+3(6-5x)=-4 5x+18-15x=-4 -10x=-22 x=2.2 y=6-5*2.2=6-11=-5 y=-5

OpenStudy (anonymous):

Thank you.